# Towards Robust Spatio-Temporal Auto-Regressive Prediction: Adams-Bashforth Time Integration with Adaptive Multi-Step Rollout

### v1

@@ 7 Comparison with conventional noise-injection approach for long-term rollout
In this section, we compare the long-term rollout performance of our models against the most widely adopted approach in the field of auto-regression: noise injection [12, 15, 16, 17]. This method deliberately adds random noise to the input data during the training phase to enhance the model’s robustness against error accumulation in AR predictions. The underlying principle is that by exposing the model to perturbed input data during training, it becomes more resilient to prediction errors that naturally accumulate during the rollout phase.
We evaluate three model configurations, all trained with Gaussian noise $\mathcal{N}(0,0.16^{2})$ injected into the input data. This noise level, identified as optimal in our previous work [17] on the same dataset, is consistently applied during training. Notably, evaluations are conducted without multi-step rollouts. The first model employs the traditional direct prediction approach, the second uses the forward Euler time integration scheme, and the third adopts the Adams-Euler time integration scheme. During inference, no additional noise is introduced, allowing the models to leverage their learned robustness to maintain performance over extended rollouts.
To assess the effectiveness of noise injection, we again analyze the temporal behavior of the $x$-velocity at seven probe points located downstream of the cylinder, as in previous evaluations (refer to Fig. 9). Figs. 10(a) and 10(b) present the results for the forward Euler and Adams-Euler scheme with noise injection, respectively. Comparing Fig. 10(a) with Fig. 9(a), which adopts forward Euler with multi-step rollout using AW2, we observe that the model with noise injection exhibits inferior accuracy at all probe points. The oscillations captured by the model deviate more significantly from the ground truth in terms of both amplitude and phase, indicating that noise injection alone is insufficient to maintain long-term prediction accuracy. Similarly, comparing Fig. 10(b) with Fig. 9(b), which adopts Adams-Euler and multi-step rollout using AW3, the model with noise injection shows less accurate reproduction of the vortex shedding patterns, especially for the lower four probe points. Although some oscillatory behavior is present, the predictions do not align closely with the ground truth, especially over extended rollout steps after 150 rollouts. Finally, it is worth noting that when comparing Figs. 10(a) and 10(b),the rollout performance of the Adams-Euler time integration scheme significantly surpasses that of the forward Euler approach, even in the case of noise injection. This observation further underscores the robust performance and superior accuracy of the Adams-Bashforth time integration scheme over the simple forward Euler method across various scenarios.
@@
(b) Adams-Bashforth with forward Euler and noise injection
@@
Table 4: MSE comparison between adaptive multi-step rollout and noise injection approaches for different time integration schemes. Each time scheme utilizes a different adaptive weighting approach for multi-step rollout, selected as the best configuration for each case. MSE values here are calculated with respect to seven probe points in Fig. 7.
@@
The performance gap between adaptive multi-step rollout and noise injection can be attributed to fundamental differences in how these methods address error accumulation. Noise injection attempts to build resilience by introducing input perturbations, making the model more robust to small deviations. However, this approach does not explicitly address the temporal dependencies and error propagation mechanisms inherent in AR predictions. In contrast, our adaptive multi-step rollout strategy enables the model to learn from multiple future steps simultaneously, using strategically adjusted loss weights to focus on both immediate and distant predictions. The effectiveness of this approach is clearly demonstrated in the time series visualizations, where models trained with adaptive multi-step rollout more accurately capture the vortex shedding patterns compared to those trained with noise injection, as shown in Figs. 9 and 10. These results establish our proposed methodology, combining Adams-Euler time integration with adaptive multi-step rollout, as a more effective and reliable alternative to the traditional noise injection technique for long-term prediction tasks.
Combination of multi-step rollout and noise injection techniques
Experiments combining multi-step rollout and noise injection were also conducted; however, their performance fell significantly short compared to methods utilizing either multi-step rollout or noise injection individually. The results suggest that the simultaneous use of these long-term rollout techniques can degrade model performance. This degradation is likely due to the excessive noise introduced into the input datasets during training, stemming both from the artificial noise added through the noise injection approach and the predictive noise generated from the multi-step rollout outputs of the AR models.
@@ 8.1 Experimental setup for truncated mesh scenario
The truncated mesh is a subset of the original computational domain, capturing the essential features of the vortex shedding phenomenon while reducing the spatial extent of the data (region of interest within the flow field is visualized in Fig. 11). We evaluate the performance of the four models that previously demonstrated satisfactory results, as reported in Table 3: these models include the direct prediction with vanilla multi-step rollout, forward Euler with adaptive weighting approach 2 (AW2), forward Euler with adaptive weighting approach 3 (AW3), and Adams-Euler combined with adaptive weighting approach 3 (AW3).
@@ 8.2 Results and discussion
Figs. 11(b)$\sim$11(e) present the predicted flow fields after 300 rollout steps from the four model settings mentioned above, alongside the ground truth in Fig. 11(a). We observe that the direct prediction (Fig. 11(b)) with vanilla multi-step rollout fails to capture the vortex shedding patterns, showing significant deviations from the ground truth. The forward Euler method with AW2 (Fig. 11(c)) shows some improvement but still struggles to accurately represent the flow structures, especially in the wake region behind the cylinder. The forward Euler method with AW3 (Fig. 11(d)) exhibits better performance than the previous two models, capturing some features of the vortex shedding but with noticeable discrepancies. In contrast, the Adams-Euler scheme combined with AW3 (Fig. 11(e)) closely matches the ground truth, successfully reproducing the vortex shedding patterns and flow dynamics within the truncated region. These results highlight the superior performance of the Adams-Bashforth with forward Euler scheme combined with adaptive weighting approach 3, even under the harsher condition of a truncated mesh.
@@
To further assess the models’ ability to capture temporal dynamics within the truncated mesh, we analyze the temporal evolution of the $x$-velocity at seven probe points strategically located in the wake region, as illustrated in Fig. 12. The results demonstrate clear performance differences across the four model configurations. The direct prediction method (Fig. 12(a), MSE: 0.019) completely fails to reproduce the periodic oscillations characteristic of vortex shedding, indicating its fundamental inability to capture flow dynamics in the truncated domain. The forward Euler method with AW2 (Fig. 12(b), MSE: 0.011) shows improved performance, successfully capturing the overall oscillatory behavior, though it gradually deviates from the ground truth in both amplitude and phase as the rollout progresses. Similarly, the forward Euler method with AW3 (Fig. 12(c), MSE: 0.013) achieves comparable accuracy to AW2, despite of the slightly lower performance. Most notably, the Adams-Euler scheme combined with AW3 (Fig. 12(d), MSE: 0.008) demonstrates superior performance, achieving the best agreement with ground truth across nearly all probe points. This configuration accurately captures the vortex shedding frequency, amplitude, and phase throughout the entire prediction horizon, with some deviations observed at probe point 4 (colored as red). These results further reinforce the robustness of the Adams-Euler scheme with adaptive multi-step rollout, even under the challenging conditions of the truncated mesh.
@@ 8.3 Final remarks on the proposed framework: Adams-Bashforth-based adaptive multi-step rollout
The evaluation under the more challenging condition of a truncated mesh provides compelling evidence for the exceptional robustness of our proposed methodology—the synergistic combination of the Adams-Bashforth with forward Euler scheme and adaptive weighting approach 3. Despite the increased complexity of predicting flow dynamics with reduced spatial information in a confined region, this integrated approach consistently demonstrates superior performance over alternative methods. While the direct prediction and forward Euler methods, even with various weighting strategies, struggle to maintain accuracy in this challenging scenario, our combined approach continues to deliver high prediction accuracy and stability. This noticeable performance contrast highlights a crucial insight: the selection of appropriate time integration schemes, when coupled with effective training strategies, plays a fundamental role in developing robust AR models.
The successful performance of our methodologies under these more challenging conditions—beyond those in our previous experiments—underscores their broad applicability and reliability. This robustness can be particularly powerful in practical engineering applications where data availability is often limited or analysis must focus on specific regions of interest within complex flow fields due to the memory limitations. By delivering accurate predictions even with truncated spatial information, our approach establishes itself as a robust and versatile tool for a wide range of real-world flow prediction scenarios. It offers a compelling solution in situations where traditional methods, such as forward Euler time integration, vanilla multi-step rollout, and noise injection, may fall short or prove insufficient.
@@ 9 Conclusions and future work
This study presents a comprehensive framework for enhancing long-term auto-regressive (AR) predictions in SciML models through the novel proposal of numerical time-integration schemes and adaptive multi-step rollout techniques. Our approach addresses the fundamental challenge of error accumulation in AR models while maintaining computational efficiency and broad applicability across various scientific domains. By combining the Adams-Bashforth forward Euler scheme with adaptive multi-step rollout strategies, we have demonstrated significant improvements in both prediction accuracy and stability over extended time horizons.
The quantitative findings obtained through the realization of each key contribution introduced in Section 1 can be summarized as follows:
  1. 1.
    Development of a comprehensive AR framework: While direct application of various time integration schemes showed minimal improvement over conventional direct prediction in long-term rollout performance, their integration with multi-step rollout revealed significant differences. The Adams-Bashforth combined with the forward Euler scheme achieved a 7% improvement in prediction accuracy compared to the commonly used forward Euler schemes, whereas the traditional direct prediction approach failed to converge. While other time integration schemes exhibited instability as the number of future time steps increases in multi-step rollout training, Adams-Bashforth with forward Euler demonstrated consistent and robust performance throughout varying prediction horizons.
  2. 2.
    Enhancement of conventional multi-step rollout through adaptive weighting: The proposed adaptive weighting strategy focusing on first and last components delivered 89% improvement over conventional fixed weights and outperformed other adaptive approaches by up to 80%, while maintaining similar computational costs. This demonstrates the effectiveness of selectively emphasizing critical time steps during training rather than considering all future steps simultaneously during multi-step rollout.
  3. 3.
    Success under challenging conditions for long-term rollout: Simulating practical scenarios where computational resources and data availability are limited, we used only 50 training snapshots and 1,177 model parameters. Under resource-constrained conditions, accurate predictions over 350 time steps were achieved by integrating the Adams-Bashforth and forward Euler schemes with adaptive weighting that prioritized the first and last loss components. This level of performance was unattainable with conventional AR techniques.
  4. 4.
    Validation of the proposed approach over traditional noise injection approach: Our approach achieved an 83% improvement in prediction accuracy compared to the traditional noise injection method, a widely used technique for enhancing the rollout performance of AR models. Besides, we observed a novel finding: combining multi-step rollout with noise injection led to degraded performance. This decline is attributed to the compounding effect of excessive noise arising from both artificial injection and errors accumulated during AR-based multi-step predictions.
  5. 5.
    Robustness in more challenging scenario with truncated mesh: In tests with a truncated mesh domain, our method achieved 58% and 27% improvement over direct prediction and forward Euler approaches respectively. This successful performance with reduced spatial information demonstrates broad applicability for practical engineering scenarios where computational memory constraints often necessitate focusing on specific narrow regions of interest within complex flow fields.
  6. 6.
    General applicability of the proposed approach: Through extensive validation, we have shown for the first time that integrating Adams-Bashforth time-stepping with adaptive multi-step rollout significantly enhances the performance of AR predictions. This versatile framework is compatible with any SciML architecture designed for AR prediction, setting a new standard for long-term temporal forecasting across diverse engineering domains.
@@
While our study marks significant advances in AR prediction capabilities, several opportunities for future research remain. First, integrating more advanced higher-order time integration schemes could further improve prediction stability and accuracy. Second, as multi-step rollout involves computing gradients across multiple future steps, developing more efficient and tactical approaches to reduce its computational overhead would be a valuable direction. Finally, extending the application of our framework to other scientific domains and exploring its performance with diverse neural network architectures and temporal dynamics would provide deeper insights into its broader utility and potential limitations, paving the way for more versatile and scalable solutions.
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (2018R1A5A7025409), and the Ministry of Science and ICT of Korea grant (No. 2022-0-00969, No. 2022-0-00986, and No. RS-2024-00355857), the Ministry of Trade, Industry & Energy grant (RS-2024-00410810), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-300). R.V. was supported by ERC Grant No. 2021-CoG-101043998, DEEPCONTROL. The views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of European Union or European Research Council.
@@ Appendix A On the applicability of Runge-Kutta methods for AR prediction
While Runge-Kutta (RK) methods, particularly the fourth-order scheme, are widely used in numerical integration due to their superior accuracy, their application to AR prediction presents fundamental challenges that make them impractical for our framework. This section explains why RK methods were not explored in our study despite their popularity in traditional numerical integration.
@@ A.2 Fundamental challenges of RK in AR context
A.2.1 Ground truth availability
The primary obstacle in implementing RK methods for AR prediction is the absence of ground truth data for intermediate stages. While $k_{1}$ can be computed directly from available data as $(\mathbf{u}(t+\Delta t)-\mathbf{u}(t))/\Delta t$, the subsequent stages $k_{2}$ and $k_{3}$ require evaluations at intermediate time points $(t+\Delta t/2)$ for which no data exists in typical spatio-temporal datasets. This presents an critical challenge for supervised training of AR models.
@@ A.2.3 Computational considerations
The RK4 method would require four separate model evaluations per time step, significantly increasing computational cost compared to our adopted Adams-Bashforth scheme. This increased cost is particularly problematic in the context of long-term predictions where hundreds of time steps may be required.

### v2

@@ 7 Comparison with conventional noise-injection approach for long-term rollout
In this section, we compare the long-term rollout performance of our models against the most widely adopted approach in the field of auto-regression: noise injection [17, 22, 23, 24]. This method deliberately adds random noise to the input data during the training phase to enhance the model’s robustness against error accumulation in AR predictions. The underlying principle is that by exposing the model to perturbed input data during training, it becomes more resilient to imperfect data by learning to handle its own flawed predictions when they are fed back as inputs during inference.
We evaluate three model configurations, all trained with Gaussian noise $\mathcal{N}(0,0.16^{2})$ injected into the input data. This noise level, identified as optimal in our previous work on the same dataset [24], is consistently applied during training. Notably, evaluations are conducted without multi-step rollouts. The first model employs the traditional direct prediction approach, the second uses the forward Euler time integration scheme, and the third adopts the Adams-Euler time integration scheme. To assess the effectiveness of noise injection, we again analyze the temporal behavior of the $x$-velocity at seven probe points located downstream of the cylinder. Figs. 11(a) and 11(b) present the results for models using noise injection with forward Euler and Adams-Euler schemes respectively. For the forward Euler integration scheme, we compare two approaches: adaptive multi-step rollout with AW2 (Fig. 10(a)) versus noise injection without multi-step rollout (Fig. 11(a)). The noise injection approach exhibits inferior accuracy at all probe points, with oscillations deviating significantly from ground truth in both amplitude and phase. This indicates that noise injection alone is insufficient to maintain long-term prediction accuracy with forward Euler integration. Similarly, for the Adams-Euler scheme, we compare AW3-based adaptive multi-step rollout (Fig. 10(b)) versus noise injection without multi-step rollout (Fig. 11(b)). The noise injection approach again shows less accurate reproduction of vortex shedding patterns, especially for the lower four probe points and after 150 rollout steps.
@@
(b) Adams-Euler with noise injection
@@
Notably, when comparing between integration schemes with noise injection (Figs. 11(a) and 11(b)), the Adams-Euler method still significantly outperforms forward Euler. This further underscores the robust performance and superior accuracy of the Adams-Bashforth time integration scheme across various training approaches.
@@
Table 4: MSE comparison between adaptive multi-step rollout and noise injection approaches for different time integration schemes. Each time scheme with multi-step rollout utilizes a different adaptive weighting approach for multi-step rollout (as denoted by footnotes), selected as the best configuration for each case. MSE values here are calculated with respect to seven probe points in Fig. 8.
@@
The performance gap between adaptive multi-step rollout and noise injection can be attributed to fundamental differences in how these methods address error accumulation. Noise injection attempts to build resilience by introducing input perturbations, making the model more robust to small deviations. However, this approach does not explicitly address the temporal dependencies and error propagation mechanisms inherent in AR predictions. In contrast, our adaptive multi-step rollout strategy enables the model to learn from multiple future steps simultaneously, using strategically adjusted loss weights to focus on both immediate and distant predictions. The effectiveness of this approach is clearly demonstrated in the time series visualizations, where models trained with adaptive multi-step rollout more accurately capture the vortex shedding patterns compared to those trained with noise injection, as shown in Figs. 10 and 11. These results establish our proposed methodology, combining Adams-Euler time integration with adaptive multi-step rollout, as a more effective and reliable alternative to the traditional noise injection technique for long-term prediction tasks.
Further investigation: integrating multi-step rollout with noise injection techniques
Experiments combining multi-step rollout and noise injection were also conducted; however, their performance fell significantly short compared to methods utilizing either multi-step rollout or noise injection individually. The results suggest that the simultaneous use of these techniques degrades model performance. This degradation occurs because multi-step rollout naturally introduces increasing prediction errors in later future steps while noise injection simultaneously adds artificial perturbations. This combination creates excessive noise that compounds throughout the prediction sequence, making the loss function unstable and preventing effective learning, particularly for distant future predictions.
@@ 8.1 Experimental setup for truncated mesh scenario
The truncated mesh is a subset of the original computational domain, capturing the essential features of the vortex shedding phenomenon while reducing the spatial extent of the data (region of interest within the flow field is visualized in Fig. 12(b)$\sim$12(e)). We evaluate the performance of the four models that previously demonstrated satisfactory results, as reported in Table 3: these models include the direct prediction with vanilla multi-step rollout, forward Euler with adaptive weighting approach 2 (AW2), forward Euler with adaptive weighting approach 3 (AW3), and Adams-Euler combined with adaptive weighting approach 3 (AW3).
@@ 8.2 Results and discussion
Figs. 12(b)$\sim$12(e) present the predicted flow fields after 300 rollout steps from the four model settings mentioned above, alongside the ground truth in Fig. 12(a). We observe that the direct prediction (Fig. 12(b)) with vanilla multi-step rollout fails to capture the vortex shedding patterns, showing significant deviations from the ground truth. The forward Euler method with AW2 (Fig. 12(c)) shows some improvement but still struggles to accurately represent the flow structures, especially in the wake region behind the cylinder. The forward Euler method with AW3 (Fig. 12(d)) and Adams-Euler scheme combined with AW3 (Fig. 12(e)) closely matche the ground truth, successfully reproducing the vortex shedding patterns and flow dynamics within the truncated region. These results highlight the superior performance of the adaptive weighting approach 3 combined with forward Euler or Adams-Euler, even under the harsher condition of a truncated mesh.
@@
To further assess the models’ ability to capture temporal dynamics within the truncated mesh, we analyze the temporal evolution of the $x$-velocity at seven probe points strategically located in the wake region, as illustrated in Fig. 13. The results demonstrate clear performance differences across the four model configurations. The direct prediction method (Fig. 13(a), MSE: 0.019) completely fails to reproduce the periodic oscillations characteristic of vortex shedding, indicating its fundamental inability to capture flow dynamics in the truncated domain. The forward Euler method with AW2 (Fig. 13(b), MSE: 0.011) shows improved performance, successfully capturing the overall oscillatory behavior, though it gradually deviates from the ground truth in both amplitude and phase as the rollout progresses. Similarly, the forward Euler method with AW3 (Fig. 13(c), MSE: 0.013) achieves comparable accuracy to AW2 in Fig. 13(b), despite of the slightly lower performance. Most notably, the Adams-Euler scheme combined with AW3 (Fig. 13(d), MSE: 0.008) demonstrates superior performance, achieving the best agreement with ground truth across nearly all probe points. This configuration accurately captures the vortex shedding frequency, amplitude, and phase throughout the entire prediction horizon, with some deviations observed at probe point 4 (colored as red). These results further reinforce the robustness of the Adams-Euler scheme with adaptive multi-step rollout, under the challenging mesh conditions.
@@ 8.3 Final remarks on the proposed framework: Adams-Bashforth-based adaptive multi-step rollout
The evaluation on a truncated mesh demonstrates the exceptional robustness of our combined Adams-Bashforth with forward Euler scheme and adaptive weighting approach 3, which consistently outperforms alternative methods despite the challenge of predicting flow dynamics with reduced spatial information. While direct prediction and forward Euler methods struggle in this scenario, our integrated approach maintains high accuracy and stability, highlighting the crucial role of appropriate time integration schemes paired with effective training strategies in developing robust autoregressive models. This performance under challenging conditions beyond our previous experiments confirms the broad applicability of our methodology, making it particularly valuable for practical engineering applications where data availability is limited or analysis must focus on specific regions within complex flow fields due to memory constraints.
@@ 9 Conclusions and future work
This study presents a comprehensive framework for enhancing long-term auto-regressive predictions in SciML models through novel numerical time-integration schemes and adaptive multi-step rollout techniques. By combining the two-step Adams-Bashforth scheme with adaptive rollout strategies, our remarkably lightweight GNN-based model—containing only 1,177 trainable parameters—demonstrates exceptional effectiveness under harsh constraints, achieving accurate predictions of complex dynamics across 350 time steps with only 1.6% error. Despite its minimal computational footprint, our methodology delivers substantial improvements: Adams-Bashforth shows 7% better accuracy over conventional methods, our adaptive strategies provide up to 89% improvement over fixed-weight approaches, and our combined approach outperforms noise injection by 83%, maintaining robustness even on truncated meshes. This powerful yet resource-efficient framework is designed to be model-agnostic, ensuring these advancements can benefit diverse scientific domains without specialized adaptations.
@@
While our study marks significant advances in AR prediction capabilities, several opportunities for future research remain. First, integrating more advanced higher-order time integration schemes could further improve prediction stability and accuracy. Second, as multi-step rollout involves computing gradients across multiple future steps, developing more efficient and tactical approaches to reduce its computational overhead would be a valuable direction. Third, incorporating physical principles from the Navier-Stokes equations directly into our framework could enhance both accuracy and physical consistency of predictions, potentially reducing the need for extensive data [32]. Finally, extending the application of our framework to other scientific domains and exploring its performance with diverse neural network architectures and temporal dynamics would provide deeper insights into its broader utility and potential limitations, paving the way for more versatile and scalable solutions.
@@ Acknowledgments
This work was supported by the National Research Foundation of Korea (2018R1A5A7025409), and the Ministry of Science and ICT of Korea grant (No. 2022-0-00969, No. 2022-0-00986, and No. RS-2024-00355857), the Ministry of Trade, Industry & Energy grant (RS-2024-00410810), and the National Research Council of Science & Technology (NST) grant by the Korea government (MSIT) (No. GTL24031-000). R.V. was supported by ERC Grant No. 2021-CoG-101043998, DEEPCONTROL. The views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of European Union or European Research Council.
@@ Appendix A On the applicability of Runge-Kutta methods for AR prediction
While Runge-Kutta (RK) methods [33, 34], particularly the fourth-order scheme, are widely used in numerical integration due to their superior accuracy, their application to AR prediction presents fundamental challenges that make them impractical for our framework. This section illustrates why RK methods were not explored in our study despite their popularity in traditional numerical integration.
@@ A.2 Fundamental challenges of RK in AR context
Although end-to-end training frameworks can, in principle, avoid the need for explicit supervision of intermediate stages, several critical issues make RK methods impractical for AR prediction in our framework.
A.2.1 Intermediate state availability
RK methods inherently rely on intermediate states (e.g., $k_{2}$ and $k_{3}$), which are evaluated at fractional time steps (e.g., $t+\Delta t/2$). In typical spatio-temporal datasets used for AR prediction, only full time-step data is available. Even if the network is trained end-to-end using only the final state, this approach undermines the original strength of the RK method. The RK4 algorithm is fundamentally based on a multi-stage time integration process that sequentially leverages previous $k$ values to accurately capture the dynamics of the system. When end-to-end training is used to mimic the RK property, the intermediate stages are not explicitly supervised or constrained by physical reasons. Consequently, the network merely combines four model outputs ($k_{1}$ through $k_{4}$) without truly reflecting the underlying multi-stage integration mechanism that is essential to RK methods. This unnecessary implementation not only fails to provide the high-accuracy benefits inherent to the RK approach but also adds complexity without following the original principles of physically motivated time integration.
@@ A.2.3 Computational considerations
While it is true that the RK4 method employs the same function $\mathbf{f}$ across its four sub-steps, the inputs and corresponding outputs for these evaluations differ significantly. Specifically, the intermediate states $k_{1}$, $k_{2}$, $k_{3}$, and $k_{4}$ are computed at different time points and with different state inputs. These differences imply that each evaluation essentially encounters a distinct distribution of inputs and produces outputs with different error characteristics. Consequently, even if one were to use a single neural network to represent $\mathbf{f}$, the network would need to generalize accurately across these varied sub-tasks—an extremely challenging requirement in practice. Therefore, the RK4 method would require four separate model evaluations per time step, significantly increasing computational cost compared to our adopted Adams-Bashforth scheme. This increased cost is particularly problematic in the context of long-term predictions where hundreds of time steps may be required.